a fair die is rolled 10 times. compute the probability that a "one" occurs on exactly one of the 10 rolls.
Answer by stanbon(75887)

a fair die is rolled 10 times. compute the probability that a "one" occurs on exactly one of the 10 rolls.
-----
For each roll P(no 1) = 9/10
For each roll P(one 1) = 1/10
----
P(x = 1) = 10C1(1/10)^1*(9/10)^9
----
P = 10*(1/10)*(9/10)^9 = 0.3874
===============================
Cheers,
Stan H.
============